LUCIFER Technical Observing, prep for Partner Science:
UT 2010 Mar 08

Observer: D.Thompson
Telescope Operator: W.Wack
Instrument Support: D.Officer
SW support: T.Edgin (on call)


  • Lost entire night to fog, wind, snow.
  • Time given to M2 team to diagnose their fiber problem

  • LBCR rear cover re-installed.
  • LBCB filter wheels checked, seem to be well aligned (just vignetting?).

  • LUCIFER Telescope Service restarted, communicates with AO2
  • Manual presets and offsets sent from LUCIFER UI, executed as expected
    • GCS triggered once, it did what it should have.

Still to do as of 10:30 MST:
  • Enter new mask data into LUCIFER engineering interface



Observing Plan

  • Closed dome SW tests, if requested
  • Expand pointto to include RDG, LDG, RFBG, MODS1
  • Correct initial collimation/pointing:
    • Use Brackett Gamma filter, can do in twilight
    • Correct IE/CA at high elevation
    • Correct IA at low elevation (AZ encoder problems)
    • Check/correct IE/CA again at high elevation
  • Open dome SW tests, if requested
  • Run dice5 pattern to check that things still work
    • Check repeated positioning on LUCIFER with and without flex comp on
  • Re-test updatePointingReference functionality.
  • Repeat measurement of focal plane scale, orientation
  • Check existing collimation model
    • Take new data if necessary
    • Stable conditions!
  • Check existing pointing model
    • Implement new collimation model if necessary
    • Stable conditions!
    • Take 12 points, with pointing log on...
    • Take full set (56 points) if the first 12 points are off.
  • Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
    • Check to make sure offsets still happen in a reasonable manner
  • Collect through-focus sequence for Andrew - reasonably dense star field.
    • also check the AGw-LUCIFER focus in cold weather for temp term?
  • Collect astrometric field data to check AGw calibration at colder temps
  • Improve map of AGw patrol field limits at high field angle
    • Requires set up work on a stone field
    • See this plot for extra patrol field area (in blue): AGw_patrolfield_v2.pdf
  • Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)
  • March 6-7, End of night - Test of improved MOS acquisition sequence

-- DavidThompson - 08 Mar 2010

-- DavidThompson - 08 Mar 2010
Topic revision: r2 - 08 Mar 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback